# The Mixed Integer Trust Region Problem

### v1

@@ 5.2 Proof of Proposition 2
$\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1)\right)$ $=\frac{2}{\epsilon^{\prime}}\operatorname{width}_{d^{\prime}}\left(\mathcal{B}_{\diamond}\right)\leq\frac{2}{\epsilon^{\prime}}p2^{p(p-1)/4}.$
@@
$\leq\frac{2(1+\epsilon^{\prime})}{\epsilon^{\prime}}p2^{p(p-1)/4}$
$=\left(2+\frac{26}{\epsilon}\right)p2^{p(p-1)/4}$
$\leq\frac{28}{\epsilon}p2^{p(p-1)/4}.$
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
$\operatorname{width}_{d}(\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime})))=\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1+\epsilon^{\prime}))\leq\frac{28}{\epsilon}p2^{p(p-1)/4}.$
@@
$=-(1+\epsilon/13)\lVert d^{\prime}\rVert_{2}+d^{\mathsf{T}}c.$
@@
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))\}\in[\rho,\rho+28p2^{p(p-1)/4}/\epsilon].$
@@
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\in[\rho,\rho+28p2^{p(p-1)/4}/\epsilon].$
@@ 6 Approximation algorithm for Equation E-MIQP
In this section we prove Theorem 1. Before diving into the proof, we present, in section Section 6.1, few additional results that we need.
@@ 6.1 Ingredients for approximation algorithm
We start by presenting two lemmas that will be instrumental in reformulating a restricted Equation E-MIQP, where the feasible region is intersected with a hyperplane, as a new Equation E-MIQP with one fewer integer variable. In the first lemma we show that, given a hyperplane $\mathcal{H}$ in $\mathbb{R}^{n}$ that satisfies some technical conditions, we can find an affine transformation that maps the hyperplane $\{y\in\mathbb{R}^{n}:y_{1}=0\}$ to $\mathcal{H}$ and that preserves mixed integer points. Lemma 13 is similar to lemma 6 in [5]. The main difference is that here we seek an affine function from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$, while in lemma 6 in [5] the affine function is from $\mathbb{R}^{n-1}$ to $\mathcal{H}$. The proof is given for completeness. It will be useful to introduce some notation for hyperplanes. A hyperplane is a set of the form
@@
where $d\in\mathbb{R}^{n}\setminus\{0\}$, $\beta\in\mathbb{R}$. We say that a hyperplane is rational, when $d\in\mathbb{Q}^{n}$ and $\beta\in\mathbb{Q}$. The arguments used in the proof are direct extensions of those for pure integer linear programs (see, e.g., [3]).
@@
Let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the set $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times n}$ unimodular, such that
@@
We define the vector $\bar{x}\in\mathbb{Z}^{n}$ by $\bar{x}_{j}:=\tilde{x}_{j}$ for $j=1,\dots,p$, and $\bar{x}_{j}:=0$ for $j=p+1,\dots,n$. We also define the unimodular matrix $T\in\mathbb{Q}^{n\times n}$ with block corresponding to the first $p$ rows and $p$ columns being equal to $U$, block corresponding to the last $n-p$ rows and $n-p$ columns being equal to the identity matrix $I_{n-p}$, and remaining entries zero, i.e.,
@@
Let $\mathcal{E}(c,Q)\subset\mathbb{R}^{n}$ be a rational ellipsoid and let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, such that
@@
From Lemma 13, there is a polynomial time algorithm that determines whether the set $\mathcal{S}$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta^{\prime}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta^{\prime}(y):=\bar{x}+T^{\prime}y$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T^{\prime}\in\mathbb{Q}^{n\times n}$ unimodular, such that
@@
$=$ $\left\{y\in\mathbb{R}^{n}:(y-({T^{\prime}}^{-1}(c-\bar{x})))^{\mathsf{T}}{T^{\prime}}^{\mathsf{T}}Q{T^{\prime}}(y-({T^{\prime}}^{-1}(c-\bar{x})))\leq 1\right\}=$
@@ 6.2 Proof of Theorem 1
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\in\left[\rho,\rho+28p2^{p(p-1)/4}/\epsilon\right].$
@@
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lceil\rho+28p2^{p(p-1)/4}/\epsilon\rceil.$
For each $\beta=\lceil\rho\rceil,\dots,\lceil\rho+28p2^{p(p-1)/4}/\epsilon\rceil$, we define the optimization problem
@@
For each $\beta$, we apply the algorithm in Lemma 14. This algorithm determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If any of the two is empty, then Equation E-MIQP-$\beta$ is infeasible, thus we now assume that both sets are nonempty. Then, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, and the preimage $\mathcal{E}^{\prime}$ of the set $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ under $\eta$. Assume first that $\mathcal{E}^{\prime}$ is a singleton $\left\{c^{\prime}\right\}$, for some $c^{\prime}\in\mathbb{Q}^{n-1}$. Then $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ is the singleton $\left\{\bar{x}+Tc^{\prime}\right\}$. We check if $\bar{x}+Tc^{\prime}\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$. If it is, then it is the optimal solution to Equation E-MIQP-$\beta$, and if it is not then Equation E-MIQP-$\beta$ is infeasible. Thus we now assume that $\mathcal{E}^{\prime}$ is an ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})\subset\mathbb{R}^{n-1}$. We then have
@@
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\mathcal{E}(c^{\prime},Q^{\prime})\\ &\quad x\in\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p},\end{split}$ (MITR-$\beta^{\prime}$)
where $\operatorname{\gamma}:=\bar{x}^{\mathsf{T}}H\bar{x}+{h}^{\mathsf{T}}\bar{x}$, $H^{\prime}:=T^{\mathsf{T}}HT$, $h^{\prime}:=T^{\mathsf{T}}{h}+2T^{\mathsf{T}}H^{\mathsf{T}}\bar{x}$. Note that Equation MITR-$\beta^{\prime}$ has the form of Equation E-MIQP, but with $p-1$ integer variables and $n-p$ continuous variables.
Step 2: Recursion. We apply recursively the algorithm to each obtained Equation MITR-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\dots,\lceil\rho+28p2^{p(p-1)/4}/\epsilon\rceil$. Since each time the problem is partitioned in Step 1, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(28p2^{p(p-1)/4}/\epsilon\right)^{p}.$
Step 3: Termination. Among all the feasible solutions to Equation E-MIQP found, we return one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from $\mathcal{E}(c,Q)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ to $\mathbb{R}$, is a continuous function from a compact set to the real numbers. Hence, from the extreme value theorem, it has a (global) minimum. ∎
@@ 7 Mixed integer quadratic programming
In this section we prove Theorem 2. Our first goal is to present a theorem of the alternative for Equation MIQP, along the lines of Proposition 2 for Equation E-MIQP: an algorithm that either finds an approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat.
7.1 Theorem of the alternative for Equation MIQP
The goal of this section is to present the following theorem of the alternative for Equation MIQP.
Proposition 3 (Theorem of the alternative for Equation MIQP).
@@ 7.1 Theorem of the alternative for Equation MIQP
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\in\left[\rho,\rho+20(n+1)p2^{p(p-1)/4}\right].$
@@
Before giving the proof, we give an overview of the algorithm. The structure is similar to the algorithm in the proof of Proposition 2, with few key differences. First, we construct an affine transformation $\tau$ that maps $\mathcal{P}$ into a polytope that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,n+1)$. We denote by Equation ${}^{\prime}$ the optimization problem obtained by applying the same affine transformation to Equation MIQP. Next, we consider Equation ${}^{\prime}$-s obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{P})$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{\#},y^{\diamond}$ to Equation ${}^{\prime}$-s such that $f(y^{\#})$ is sufficiently larger than $y^{\diamond}$. If Proposition 1 finds the vectors $y^{\#},y^{\diamond}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation ${}^{\prime}$, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. Otherwise, we show that $\mathcal{P}$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 3, which uses Lemmas 10, 1 and 12.
@@
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{P})\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})-\{Bc\},\end{split}$ (${}^{\prime}$)
where $\operatorname{\gamma}:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, and $b^{1},\dots,b^{n}$ are the columns of $B$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider now the mixed integer trust region problem obtained from Equation ${}^{\prime}$ by replacing the constraint $y\in\tau(\mathcal{P})$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\operatorname{\gamma}+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})-\{Bc\}.\end{split}$ (${}^{\prime}$-s)
Using the algorithm in Proposition 1 (with $\epsilon:=1/10$), we either find two feasible solutions $y^{\#},y^{\diamond}$ to Equation ${}^{\prime}$-s such that
@@
In the remainder of the proof, we consider separately two cases. First, consider the case where Proposition 1 successfully found the two vectors $y^{\#},y^{\diamond}$. Note that $y^{\#}$ and $y^{\diamond}$ are feasible also to Equation ${}^{\prime}$. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation ${}^{\prime}$. From (7), we directly obtain the following lower bound on $f_{\sup}-f(y^{\diamond})$:
@@
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation ${}^{\prime}$, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. We combine (8) and (9) to obtain:
$f_{\sup}-f_{\inf}\leq 4(n+1)^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\leq 8(n+1)^{2}\left(f_{\sup}-f(x^{\diamond})\right).$
@@
This concludes the first case.
@@
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,n+1))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
@@
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,n+1))\}\in[\rho,\rho+20(n+1)p2^{p(p-1)/4}].$

### v2

@@ 6.2 Proof of Proposition 2
$\operatorname{width}_{d^{\prime}}\left(\mathcal{B}(0,1)\right)$ $=\frac{1}{\epsilon^{\prime}}\operatorname{width}_{d^{\prime}}\left(\mathcal{B}_{\diamond}\right)\leq\frac{1}{\epsilon^{\prime}}p2^{p(p-1)/4}.$
@@
$\leq\frac{(1+\epsilon^{\prime})}{\epsilon^{\prime}}p2^{p(p-1)/4}$
$=\left(1+\frac{18}{\epsilon}\right)p2^{p(p-1)/4}$
$\leq\frac{19}{\epsilon}p2^{p(p-1)/4}.$
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
$\operatorname{width}_{d}(\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime})))=\operatorname{width}_{d^{\prime}}(\mathcal{B}(0,1+\epsilon^{\prime}))\leq\frac{19}{\epsilon}p2^{p(p-1)/4}.$
@@
$=-(1+\epsilon/18)\lVert d^{\prime}\rVert_{2}+d^{\mathsf{T}}c.$
@@
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,1+\epsilon^{\prime}))\}\subseteq[\rho,\rho+19p2^{p(p-1)/4}/\epsilon].$
@@
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\subseteq[\rho,\rho+19p2^{p(p-1)/4}/\epsilon].$
@@ 7 Approximation algorithm for Equation E-MIQP
In this section we prove Theorem 1. Before diving into the proof, we present, in Section 7.1, a few additional results that we need.
@@ 7.1 Ingredients for approximation algorithm
We start by presenting two lemmas that will be instrumental in reformulating a restricted Equation E-MIQP, where the feasible region is intersected with a hyperplane, as a new Equation E-MIQP with one fewer integer variable. In the first lemma we show that, given a hyperplane $\mathcal{H}$ in $\mathbb{R}^{n}$ that satisfies some technical conditions, we can find an affine transformation that maps the hyperplane $\{y\in\mathbb{R}^{n}:y_{1}=0\}$ to $\mathcal{H}$ and that preserves mixed integer points. Lemma 13 is similar to lemma 6 in [7]. The main difference is that here we seek an affine function from $\mathbb{R}^{n}$ to $\mathbb{R}^{n}$, while in lemma 6 in [7] the affine function is from $\mathbb{R}^{n-1}$ to $\mathcal{H}$. It will be useful to introduce some notation for hyperplanes. A hyperplane is a set of the form
@@
where $d\in\mathbb{R}^{n}\setminus\{0\}$, $\beta\in\mathbb{R}$. We say that a hyperplane is rational, when $d\in\mathbb{Q}^{n}$ and $\beta\in\mathbb{Q}$. Recall that a square matrix $T\in\mathbb{Z}^{n\times n}$ is called unimodular if it has determinant ±$\pm 1$. The arguments used in the proof are direct extensions of those for pure integer linear programs (see, e.g., [4]).
@@
Let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the set $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Z}^{n\times n}$ unimodular, such that
@@
We define the vector $\bar{x}\in\mathbb{Z}^{n}$ by $\bar{x}_{j}:=\tilde{x}_{j}$ for $j=1,\dots,p$, and $\bar{x}_{j}:=0$ for $j=p+1,\dots,n$. We also define the unimodular matrix $T\in\mathbb{Z}^{n\times n}$ with block corresponding to the first $p$ rows and $p$ columns being equal to $U$, block corresponding to the last $n-p$ rows and $n-p$ columns being equal to the identity matrix $I_{n-p}$, and remaining entries zero, i.e.,
@@
Let $\mathcal{E}(c,Q)\subset\mathbb{R}^{n}$ be a rational ellipsoid and let $\mathcal{H}(d,\beta)\subset\mathbb{R}^{n}$ be a rational hyperplane with $d_{p+1}=\cdots=d_{n}=0$, for $p\in\{1,\dots,n\}$. There is a polynomial time algorithm which determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If $\mathcal{S}\neq\emptyset$ and $n\geq 2$, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, such that
@@
From Lemma 13, there is a polynomial time algorithm that determines whether the set $\mathcal{S}$ is empty or not. If $\mathcal{S}\neq\emptyset$, the algorithm finds a map $\eta^{\prime}:\mathbb{R}^{n}\to\mathbb{R}^{n}$ of the form $\eta^{\prime}(y):=\bar{x}+T^{\prime}y$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T^{\prime}\in\mathbb{Z}^{n\times n}$ unimodular, such that
@@
$=$ $\left\{y\in\mathbb{R}^{n}:(y-({T^{\prime}}^{-1}(c-\bar{x})))^{\mathsf{T}}{T^{\prime}}^{\mathsf{T}}Q{T^{\prime}}(y-({T^{\prime}}^{-1}(c-\bar{x})))\leq 1\right\}$
@@ 7.2 Proof of Theorem 1
$\{d^{\mathsf{T}}x:x\in\mathcal{E}(c,Q)\}\subseteq\left[\rho,\rho+19p2^{p(p-1)/4}/\epsilon\right].$
@@
$\{x\in\mathbb{R}^{n}:{d}^{\mathsf{T}}x=\beta\},\qquad\beta=\lceil\rho\rceil,\ \lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor.$
For each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor$, we define the optimization problem
@@
If $n=1$, Equation E-MIQP-$\beta$ can be trivially solved, thus we now assume $n\geq 2$. For each $\beta$, we apply the algorithm in Lemma 14. This algorithm determines whether the sets $\mathcal{S}:=\mathcal{H}(d,\beta)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ and $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ are empty or not. If any of the two is empty, then Equation E-MIQP-$\beta$ is infeasible, thus we now assume that both sets are nonempty. Then, the algorithm finds a map $\eta:\mathbb{R}^{n-1}\to\mathbb{R}^{n}$ of the form $\eta(y):=\bar{x}+Ty$, with $\bar{x}\in\mathbb{Z}^{n}$ and $T\in\mathbb{Q}^{n\times{n-1}}$ of full rank, and the preimage $\mathcal{E}^{\prime}$ of the set $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ under $\eta$. Assume first that $\mathcal{E}^{\prime}$ is a singleton $\left\{c^{\prime}\right\}$, for some $c^{\prime}\in\mathbb{Q}^{n-1}$. Then $\mathcal{E}(c,Q)\cap\mathcal{H}(d,\beta)$ is the singleton $\left\{\bar{x}+Tc^{\prime}\right\}$. We check if $\bar{x}+Tc^{\prime}\in\mathbb{Z}^{p}\times\mathbb{R}^{n-p}$. If it is, then it is the optimal solution to Equation E-MIQP-$\beta$, and if it is not then Equation E-MIQP-$\beta$ is infeasible. Thus we now assume that $\mathcal{E}^{\prime}$ is an ellipsoid $\mathcal{E}(c^{\prime},Q^{\prime})\subset\mathbb{R}^{n-1}$. We then have
@@
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\mathcal{E}(c^{\prime},Q^{\prime})\\ &\quad x\in\mathbb{Z}^{p-1}\times\mathbb{R}^{n-p},\end{split}$ (MITR-$\beta^{\prime}$)
where $\gamma:=\bar{x}^{\mathsf{T}}H\bar{x}+{h}^{\mathsf{T}}\bar{x}$, $H^{\prime}:=T^{\mathsf{T}}HT$, $h^{\prime}:=T^{\mathsf{T}}{h}+2T^{\mathsf{T}}H^{\mathsf{T}}\bar{x}$. Note that Equation MITR-$\beta^{\prime}$ has the form of Equation E-MIQP, but with $p-1$ integer variables and $n-p$ continuous variables.
Step 2: Recursion. We apply recursively the algorithm to each obtained Equation MITR-$\beta^{\prime}$, for each $\beta=\lceil\rho\rceil,\lceil\rho\rceil+1,\dots,\lfloor\rho+19p2^{p(p-1)/4}/\epsilon\rfloor$. Since each time the problem is partitioned in Step 1, the number of integer variables decreases by one, the total number of iterations of the algorithm is upper bounded by
$\left(19p2^{p(p-1)/4}/\epsilon\right)^{p}.$
Step 3: Termination. If the algorithm did not find any feasible solution to Equation E-MIQP, the algorithm correctly returns that the feasible region is empty. Otherwise, among all the feasible solutions to Equation E-MIQP found, the algorithm returns one with the minimum objective value. The result then follows from Lemma 15. This is because the function $f(x):=x^{\mathsf{T}}Hx+h^{\mathsf{T}}x$, from $\mathcal{E}(c,Q)\cap(\mathbb{Z}^{p}\times\mathbb{R}^{n-p})$ to $\mathbb{R}$, is a continuous function from a compact set to the real numbers. Hence, from the extreme value theorem, it has a (global) minimum. ∎
@@ 8 Mixed integer quadratic programming
In this section we prove Theorem 2.
8.1 Flatness result for Equation MIQP
Our first goal is to present a flatness result for Equation MIQP: an algorithm that either finds an approximate solution, or finds a nonzero vector $d$ along which the feasible region is flat. This result has several similarities with the flatness result for Equation E-MIQP that we presented in Proposition 2. The key difference, besides the optimization problem considered, is that the algorithm in Proposition 2 can be applied with any $\epsilon\in(0,1]$, while the algorithm in the flatness result for Equation MIQP only works with $\epsilon=1-\Theta(n^{-2})$.
Proposition 3 (Flatness result for Equation MIQP).
@@ 8.1 Flatness result for Equation MIQP
$\left\{d^{\mathsf{T}}x:Wx\leq w\right\}\subseteq\left[\rho,\rho+14(n+1)p2^{p(p-1)/4}\right].$
@@
Before giving the proof, we give an overview of the algorithm. The structure is similar to the algorithm in the proof of Proposition 2, with few key differences. First, we construct an affine transformation $\tau$ that maps $\mathcal{P}$ into a polytope that is sandwiched between $\mathcal{B}(0,1)$ and $\mathcal{B}(0,n+1)$. We denote by Equation MIQP the optimization problem obtained by applying the same affine transformation to Equation MIQP. Next, we consider Equation MIQP-s obtained from Equation MIQP by replacing the constraint $y\in\tau(\mathcal{P})$ with with the stronger requirement $y\in\mathcal{B}(0,1)$. Using Proposition 1, we search for two feasible solutions $y^{\#},y^{\diamond}$ to Equation MIQP-s such that $f(y^{\#})$ is sufficiently larger than $y^{\diamond}$. If Proposition 1 finds the vectors $y^{\#},y^{\diamond}$, we show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. Otherwise, we show that $\mathcal{P}$ is flat and we find the associated vector $d$. We now give the complete proof of Proposition 3, which uses Lemmas 10, 1 and 12.
@@
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y\in\tau(\mathcal{P})\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\},\end{split}$ (MIQP)
where $\gamma:=c^{\mathsf{T}}Hc+{h}^{\mathsf{T}}c$, $H^{\prime}:=B^{-\mathsf{T}}HB^{-1}$, $h^{\prime}:=2B^{-\mathsf{T}}H^{\mathsf{T}}c+B^{-\mathsf{T}}h$, $b^{1},\dots,b^{n}$ are the columns of $B$, and $c^{\prime}:=-Bc$. For ease of notation, in the following we denote by $f_{y}(y):=y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y$.
Consider now the mixed integer trust region problem obtained from Equation MIQP by replacing the constraint $y\in\tau(\mathcal{P})$ with the stronger requirement $y\in\mathcal{B}(0,1)$:
$\begin{split}\gamma+\min&\quad y^{\mathsf{T}}H^{\prime}y+{h^{\prime}}^{\mathsf{T}}y\\ \operatorname{s.t.}&\quad y^{\mathsf{T}}y\leq 1\\ &\quad y\in\Pi_{p}(b^{1},\dots,b^{n})+\{c^{\prime}\}.\end{split}$ (MIQP-s)
Using the algorithm in Proposition 1 (with $\epsilon:=1/14$), we either find two feasible solutions $y^{\#},y^{\diamond}$ to Equation MIQP-s such that
@@
In the remainder of the proof, we consider separately two cases. First, consider the case where Proposition 1 successfully found the two vectors $y^{\#},y^{\diamond}$. Note that $y^{\#}$ and $y^{\diamond}$ are feasible also to Equation MIQP. Denote by $f_{\sup}$ and $f_{\inf}$ the supremum and the infimum of $f_{y}(y)$ on the feasible region of Equation MIQP. From (9), we directly obtain the following lower bound on $f_{\sup}-f(y^{\diamond})$:
@@
We are now ready to show that $y^{\diamond}$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. We combine (10) and (11) to obtain:
$f_{\sup}-f_{\inf}\leq 4(n+1)^{2}\max\left\{\lVert H^{\prime}\rVert_{2},\lVert h^{\prime}\rVert_{2}\right\}\leq 8(n+1)^{2}\left(f_{\sup}-f(y^{\diamond})\right).$
@@
It then follows that $x^{\diamond}:=B^{-1}y^{\diamond}+c$ is an $\epsilon$-approximate solution to Equation MIQP, where the approximation factor is $\epsilon=1-\Theta(n^{-2})$. This concludes the first case.
@@
Now let $d:=B^{\mathsf{T}}{d^{\prime}}\in\mathbb{Q}^{n}$. From Lemma 12, $d\in\mathbb{Z}^{n}$, $d_{p+1}=\cdots=d_{n}=0$, and, for any nonempty bounded closed set $\mathcal{S}\subseteq\mathbb{R}^{n}$, $\operatorname{width}_{d}(\mathcal{S})=\operatorname{width}_{d^{\prime}}(\tau(\mathcal{S})).$ If we let $\mathcal{S}:=\tau^{\leftarrow}(\mathcal{B}(0,n+1))$, where $\tau^{\leftarrow}$ denotes the inverse of $\tau$, we obtain
@@
$\{d^{\mathsf{T}}x:x\in\tau^{\leftarrow}(\mathcal{B}(0,n+1))\}\subseteq[\rho,\rho+14(n+1)p2^{p(p-1)/4}].$